What is the mass in grams of  1.10 x 10^24 atoms of lithium?

......grams


How many lithium atoms are there in 11.2 grams of lithium?

......atoms


1
Expert's answer
2021-03-01T03:24:44-0500

(i)

Number of moles of Li atom"=\dfrac{1.10\times 10^{24}}{6.022\times 10^{23}}=1.826mol"


Mass of Lithium="\text{ no. of moles }\times \text{molar mass}=1.826\times 6.941=12.67g"




(ii)

Given, Mass of Lithim=11.2g


Moles of Lithium in "11.2g=\dfrac{11.2}{6.941}=1.6136"


Number of atoms "=\text{Number of moles }\times \text{Avagadro number}"


"=1.6136\times 6.022\times 10^{23}=9.717\times 10^{23} \text{atoms}"